# GARCH copulas and GARCH-mimicking copulas

### v1

@@ 6.2 Tractable processes using linear v-transforms
Equation (28) describes a vt-d-vine copula process. In such a model the copula density can be evaluated by first applying the v-transform to the entire vector $(u_{1},\ldots,u_{n})$ and then passing the resulting values to the density of a d-vine copula based on the sequence $(C_{k}^{*})_{k\in\mathbb{N}}$. McNeil, (2021) shows that if we apply such a copula process to a series $(X_{t})_{t\in\mathbb{Z}}$ we effectively use the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$ to model the serial dependence structure of an element-wise transformation of the time series $(T(X_{t}))_{t\in\mathbb{Z}}$. The case where the copulas are all Gaussian corresponds to the vt-ARMA model in McNeil, (2021), although this tends to be an inferior model to the ones based on non-Gaussian d-vines in Bladt and McNeil, 2022b . The implied canonical form of $T$ for a linear v-transform $\mathcal{V}(\cdot;\delta)$ and a continuous marginal distribution function $F_{X}$ is the v-shaped function on $\mathbb{R}$ given by
$T(x)=\begin{cases}F_{X}^{-1}\left(F_{X}(x)+\mathcal{V}(F_{X}(x);\delta)\right)-\mu&x\leqslant\mu,\\ x-\mu&x>\mu,\end{cases}$ (30)
where $\mu=F_{X}^{-1}(\delta)$ [^5] The function $T$ in (30) is not the unique transformation that gives a time series whose dependence structure may be described by the joint density (28). There is, in fact, an equivalence class of transformations that all lead to the same copula density when paired with $F_{X}$. For example, in the case where $F_{X}$ is symmetric and $\delta=0.5$ we could also take $T(x)=x^{2}$. We pin down the canonical form by fixing the gradient to be 1 for $x>\mu$. See McNeil, (2021) for more details.. For example, in the case where $F_{X}$ is symmetric around the origin and $\delta=0.5$, $\mu=0$ and $T(x)=|x|$. In general $T$ can be thought of as a function that measures the magnitude of a return relative to $\mu$ in a way that can also account for asymmetric effects for returns that are greater or less than $\mu$.
7 Conclusion
This paper has asked two questions. First, what are the bivariate copulas that underlie the serial dependencies and serial conditional dependencies of pairs of variables $(X_{t},X_{t+k})$ in first-order GARCH(1,1) models. Second, can these copulas be mimicked by simpler constructions with a view to developing simplified d-vines that can behave in a similar fashion to GARCH-type models, but offer more flexible solutions for datasets where GARCH works less well.
With regard to the first question we find interesting symmetries in the dependence structures of the models with symmetric innovation distributions, even when leverage effects are introduced, but these symmetries disappear in the presence of asymmetric innovations, which are often necessary in practice to model real-world datasets. What can generally always be said is that the underlying bivariate copulas are cross-shaped, in the sense that there is a tendency for points to accumulate in all fours corners of the unit square, and that this tendency increases at shorter lags as the size of the ARCH effect increases and persists at longer lags as the aggregate size of the ARCH and GARCH effects increases. Heavier-tailed innovation distributions lead to weaker dependence.
However, the true copulas and conditional copulas of GARCH-type processes remain inaccessible for practical use, mainly because the stationary distribution of these models is not available in closed form. This leads us to consider the second question of whether GARCH copulas can be mimicked with simpler copulas. Motivated by the recognition that v-transforms are implicitly embedded in the copulas of symmetric GARCH models, we propose a construction based on v-transforms and positive-dependence copulas with strong dependence in the upper joint tail; these constructions are naturally cross-shaped. Of the well-known copulas, Joe’s copula in combination with asymmetric v-transforms seems the most promising for mimicking GARCH. However, we also identify a new copula - the absolute spherical t copula - which seems to offer an even more promising basis for the construction.
Could these copulas be used to construct simplified d-vines to mimic GARCH processes? To this question we give a more qualified answer. While the Markovian first-order ARCH-type models can definitely be mimicked by a d-vine(1) models, as soon as a GARCH effect is introduced, it is not clear whether the simplifying assumption will restrict our ability to perfectly mimic GARCH behaviour; a first investigation of this question in Section 5 produces some weak evidence that the dependence of the conditional copulas on the conditioning variables may not be particularly strong, but more investigation is required. Moreover, the computational requirements of d-vines, as discussed in Section 6, mean that we may have difficulty in integrating the most promising GARCH-mimicking pair copulas (such as iv2-iv2-ast) into d-vine structures and we may have to compromise and use copulas based on linear v-transforms. Further examination of these issues is required.
Could the GARCH-mimicking copulas be used to construct d-vines to model real financial return data? This is a different question to the question of mimicking GARCH processes. Bladt and McNeil, 2022b  have shown that the class of vt-d-vine processes with linear v-transforms is competitive for modelling many financial return series and, in particular, can give better results for data which depart from the power-lawed marginal behaviour modelled by a GARCH process. The d-vine processes constructed from GARCH-mimicking pair copulas, as described in this paper, can offer additional flexibility because of the use of two v-transforms in their construction, even if we do restrict both to be linear. There is a need for further empirical research to evaluate the potential of such models.
Another issue requiring further research is the construction of d-vines to mimic ARMA-GARCH processes. Real-world financial return series sometimes show serial correlation in the raw data as well as in the squared or absolute value series and this is usually modelled by combining ARMA and GARCH effects. In this paper we have concentrated on pure GARCH-type models and given only a single example in Section 3.4 of an AR-ARCH model. The illustration in the lower-right panel of Figure 3 suggests mixtures of positive dependence copulas and GARCH-mimicking copulas could be successful in addressing combined ARMA and GARCH effects. As a pragmatic alternative, the d-vine models discussed in this paper could be applied to the residuals from an ARMA model instead.

### v2

@@ 5 Conclusion
Equation (S9) describes a vt-D-vine copula process. In such a model the copula density can be evaluated by first applying the v-transform to the entire vector $(u_{1},\ldots,u_{n})$ and then passing the resulting values to the density of a d-vine copula based on the sequence $(C_{k}^{*})_{k\in\mathbb{N}}$. McNeil, (2021) shows that if we apply such a copula process to a series $(X_{t})_{t\in\mathbb{Z}}$ we effectively use the copula sequence $(C_{k}^{*})_{k\in\mathbb{N}}$ to model the serial dependence structure of the time series $(T(X_{t}))_{t\in\mathbb{Z}}$ where $T$ is a v-shaped transformation on $\mathbb{R}$.
Note that the Theorem above is a little more general than is needed. The copula sequence in our application is derived from a single parametric copula family $C^{*}$ and we have $C_{k}^{*}(u,v)=C^{*}(u,v;\theta_{k})$ for a sequence of parameters $(\theta_{k})_{k\in\mathbb{N}}$. However, we could obviously choose the $C_{k}^{*}$ from different parametric families.
Appendix S7 Detailed estimation results
Table S3 contains the estimation results for various D-vine models fitted to 10000 simulated data from a GARCH(1,1) process with symmetric innovations from a t distribution with 4 degrees of freedom. The most effective models in terms of AIC and BIC are the ast-AR(5) and ast-ARMA(1,1) models.
Model D-vine order No. pars Loglik AIC BIC
T1 1 2 442.8 -881.6 -867.2
B1 1 5 436.6 -863.2 -827.1
Joe-AR(1) 1 3 440.6 -875.1 -853.5
Clayton180-AR(1) 1 3 424.6 -843.2 -821.6
ast-AR(1) 1 3 442.8 -879.5 -857.9
T2 2 4 576.3 -1144.6 -1115.7
B2 2 10 568.6 -1117.3 -1045.2
Joe-AR(2) 2 4 564.4 -1120.7 -1091.9
Clayton180-AR(2) 2 4 552.1 -1096.2 -1067.4
ast-AR(2) 2 4 573.1 -1138.2 -1109.4
T5 5 10 695.6 -1371.3 -1299.2
B5 5 25 689.7 -1329.4 -1149.2
Joe-AR(5) 5 7 681.9 -1349.8 -1299.3
Clayton180-AR(5) 5 7 667.9 -1321.9 -1271.4
ast-AR(5) 5 7 693.9 -1373.7 -1323.3
Joe-ARMA(1,1) 20 4 682.0 -1356.0 -1327.2
Clayton180-ARMA(1,1) 20 4 669.5 -1331.0 -1302.2
ast-ARMA(1,1) 20 4 694.0 -1380.0 -1351.1
Table S3: Comparison of AIC and BIC values when various D-vines are fitted to simulated data from a GARCH(1,1) process with symmetric t4 innovations.
Table S4 contains corresponding results for 10000 data from a GARCH(1,1) process with skewed t4 innovations while Table S5 is for a model which has both skewed t4 innovations and leverage. Recall that parameter values may be found in Table 1. In both these analyses the preferred models are ast-AR(5) and ast-ARMA(1,1). The improvement of ast-ARMA(1,1) on ast-AR(5) is not quite so large as in the case of the empirical data in Table 4, since the persistence of volatility in the simulated GARCH processes is lower.
Model D-vine order No. pars Loglik AIC BIC
T1 1 2 503.0 -1002.1 -987.7
B1 1 5 499.5 -989.0 -953.0
Joe-AR(1) 1 3 500.8 -995.6 -974.0
Clayton180-AR(1) 1 3 481.1 -956.2 -934.6
ast-AR(1) 1 3 505.5 -1005.1 -983.4
T2 2 4 649.2 -1290.5 -1261.7
B2 2 10 650.7 -1281.4 -1209.3
Joe-AR(2) 2 4 648.5 -1289.0 -1260.1
Clayton180-AR(2) 2 4 626.3 -1244.6 -1215.7
ast-AR(2) 2 4 656.4 -1304.9 -1276.0
T5 5 10 747.5 -1475.0 -1402.9
B5 5 25 756.4 -1462.8 -1282.6
Joe-AR(5) 5 7 744.4 -1474.7 -1424.2
Clayton180-AR(5) 5 7 723.0 -1431.9 -1381.4
ast-AR(5) 5 7 758.4 -1502.8 -1452.4
Joe-ARMA(1,1) 20 4 744.5 -1481.0 -1452.2
Clayton180-ARMA(1,1) 20 4 726.5 -1445.0 -1416.1
ast-ARMA(1,1) 20 4 759.1 -1510.3 -1481.4
Table S4: Comparison of AIC and BIC values when various D-vines are fitted to simulated data from a GARCH(1,1) process with skew t4 innovations.
Model D-vine order No. pars Loglik AIC BIC
T1 1 2 569.2 -1134.4 -1120.0
B1 1 5 588.6 -1167.3 -1131.2
Joe-AR(1) 1 3 590.2 -1174.3 -1152.7
Clayton180-AR(1) 1 3 568.5 -1130.9 -1109.3
ast-AR(1) 1 3 597.6 -1189.1 -1167.5
T2 2 4 723.8 -1439.5 -1410.7
B2 2 10 760.5 -1501.0 -1428.8
Joe-AR(2) 2 4 754.4 -1500.8 -1472.0
Clayton180-AR(2) 2 4 725.6 -1443.3 -1414.4
ast-AR(2) 2 4 763.6 -1519.1 -1490.3
T5 5 10 813.8 -1607.6 -1535.5
B5 5 25 888.2 -1726.5 -1546.3
Joe-AR(5) 5 7 868.1 -1722.1 -1671.6
Clayton180-AR(5) 5 7 828.8 -1643.6 -1593.1
ast-AR(5) 5 7 879.6 -1745.1 -1694.6
Joe-ARMA(1,1 20 4 866.0 -1724.0 -1695.1
Clayton180-ARMA(1,1) 20 4 831.0 -1654.1 -1625.2
ast-ARMA(1,1) 20 4 880.0 -1752.0 -1723.2
Table S5: Comparison of AIC and BIC values when various D-vines are fitted to simulated data from a GARCH(1,1) process with skew t4 innovations and leverage.
Appendix S8 Analysis of VaR exceedances
After model estimation, the Rosenblatt functions $R_{k}$ in (S6) can be used to construct one-step-ahead conditional quantile or value-at-risk (VaR) forecasts for all our fitted D-vine copula models. We follow Loaiza-Maya et al., (2018) by using this as the basis of a method for evaluating the quality of the fitted models. Specifically, for $t=2,\ldots,n=3668$ we compare the realized values $u_{t}$ with VaR forecasts $\operatorname{VaR}_{t}(\alpha)$ derived according to the formula
$\operatorname{VaR}_{t}(\alpha)=R_{k(t)}^{-1}(\alpha;u_{t-1},\ldots,u_{t-k(t)})$
for $\alpha\in\{0.01,0.05,0.1,0.9,0.95,0.99\}$. The number of conditioning variable in the forecast is taken to be $k(t)=\min(t-1,p,12)$ where $p$ is the Markov order of the fitted model and 12 is a cap on the Markov order to ensure reasonably fast run times. Note that, we do not compare the realized raw data $x_{t}$ with estimates of the conditional quantiles of the composite model for $(X_{t})$, but rather we compare the data transformed to the copula scale $u_{t}=\widehat{F}_{X}(x_{t})$ with estimates of the conditional quantiles of the D-vine process $(U_{t})$; this means that we evaluate the quality of the copula model independently of the quality of the adaptive kernel density estimator $\widehat{F}_{X}$ of Loaiza-Maya et al., (2018) which is used for the marginal distribution $F_{X}$.
The hit rates $\hat{\alpha}=\tfrac{1}{n-1}\sum_{t=2}^{n}I_{\{u_{t}<\operatorname{VaR}_{t}(\alpha)\}}$ for our favoured models (Joe-AR(5), ast-AR(5), Clayton180-ARMA(1,1) and ast-ARMA(1,1)) are shown in Table S6 and are very similar to the rates reported by Loaiza-Maya et al., (2018). In all cases the conditional coverage test of Christoffersen, (1998) leads to an insignificant test result with confidence level 99%.
Quantile level $\alpha$ (percent)
Model 1 5 10 90 95 99
Joe-AR(5) 0.98 4.85 9.54 90.35 95.07 99.13
ast-AR(5) 0.95 5.04 9.60 90.08 94.98 99.18
ast-ARMA(1,1) 1.06 4.83 9.87 90.21 94.98 99.05
Clayton180-ARMA(1,1) 1.04 4.88 10.33 90.19 94.77 99.15
Table S6: Empirical hit rates (expressed as percentages) for 1-day-ahead conditional $\alpha$-quantile (VaR) estimates in analysis of USD/AUD exchange-rate returns. For details of models see Section 4.2 and Table 4.
Appendix S9 Copula of an AR(1) model with ARCH(1) errors
In this section we consider a time series model with stochastic volatility as well as a non-zero stochastic conditional mean term. Specifically we consider an AR(1) model with ARCH(1) errors which is a first-order Markov process $(X_{t})_{t\in\mathbb{Z}}$ satisfying equations
$X_{t}=\phi X_{t-1}+\sigma(X_{t-1})\;\epsilon_{t}\quad\text{with}\quad\sigma(x)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}},$ (S11)
where the innovations $(\epsilon_{t})_{t\in\mathbb{Z}}$ have a symmetric distribution. This model was proposed and analysed by Borkovec and Klüppelberg, (2001) who showed that it is covariance stationary provided $\phi^{2}+\alpha_{1}<1$ and has a symmetric stationary distribution $F_{X}$. Note that it differs from the usual version of the AR(1)-ARCH(1) model, $X_{t}=\phi X_{t-1}+\sigma(X_{t-1}-\phi X_{t-2})\;\epsilon_{t}$, but this is not first-order Markov and thus less straightforward to treat with our methods.
It may be easily verified that the copula density satisfies c(u,v)=c(1u,1v)c(u,v)=c(1-u,1-v) and so the copula is radially symmetric. However, the copula is neither h-symmetric nor v-symmetric, and thus it is not jointly symmetric. The copula density is shown in the bottom right panel of Figure S2 in the case where $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$. It is evident that more of the probability mass is now on the main diagonal of the copula. Behaviour of this kind could perhaps be modelled by a mixture of a Gauss copula (or Frank copula) and one of the inverse-v-transformed copulas examined earlier.